OpenStudy (vuriffy):

It is given that:

OpenStudy (vuriffy):

\[\int\limits_{0}^{a} (\frac{ 1 }{ 2 }e^{3x} +x^2) dx = 10 \]

OpenStudy (vuriffy):

Show that \[a = \frac{ 1 }{ 3 }\ln(61-2a^3)\]

OpenStudy (vuriffy):

@518nad

OpenStudy (518nad):

okay

OpenStudy (518nad):

integration is linear so u can integrate one term at a time

OpenStudy (vuriffy):

Okay, one second.

OpenStudy (vuriffy):

I forgot how to do the e part, actually.

OpenStudy (518nad):

okay so remember e^x is sepcial because d/dx e^x = e^x and therefore int e^x dx = e^x

OpenStudy (518nad):

d/dx e^3x = (3x)' e^(3x) = 3*e^3x by chain rule, so you know that integral of 3*e^3x = e^3x very close to what we got this means integral of e^3x = 1/3 e^3x

OpenStudy (vuriffy):

So it should be (e^3x)/6? or 1/6 e^3x (same thing)

OpenStudy (518nad):

yes

OpenStudy (vuriffy):

\[\int\limits_{?}^{?} (e ^{3x})/(6) + \int\limits_{?}^{?} (x^3)/(3)\]

OpenStudy (518nad):